OpenStudy (anonymous):

33x27= use the product of a sum and a difference to evaluate the expression

OpenStudy (mayankdevnani):

(30+3)(30-3)

OpenStudy (mayankdevnani):

30^3-3^3 900-9=891

OpenStudy (mayankdevnani):

ok @cdelomas

OpenStudy (anonymous):

thanks

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{33 \times 27=(30+3)(30-3)}}\]\[\Huge{\color{green}{(30)^2-(3)^2=900-9}}\]\[\Huge{\color{blue}{891}}\]